OpenStudy (anonymous):

how to solve 154a4b3c2/14ab3c2. It wants it to write the fraction in simplest form.

OpenStudy (watchmath):

\(\frac{14ab^3c^2\cdot 11a^3}{14ab^3c^2\cdot }=11a^3\)

OpenStudy (anonymous):

thats wat i put and i got it wrong

OpenStudy (anonymous):

thanks i see wat i did wrong